# Can Emulating Semantic Translation Help LLMs with Code Translation? A Study Based on Pseudocode

### v2

@@ 4.5. Discussion: Advantages and Limitations of Pseudocode-Based Translation
After understanding the effectiveness of pseudocode-based code translation from the quantitative results in four RQs, we further conduct case studies to attribute its successes and failures to learn concrete insights for the effective adoption of this approach. Based on these findings, we also suggest several research directions to mitigate the limitations and harness the potential.
Key advantages of pseudocode-based code translation. We identified three major advantages of emulating explicit semantic translation via pseudocode-based translation to complement direct translation. One of them facilitates the understanding of the original programs and reduces the burden of one step, and the other two help LLMs handle differences across PLs. Practitioners are encouraged to leverage the advantages of pseudocode-based translation for original programs or PL pair fitting them.
@@ 1. Pro-1
    Pseudocode helps disentangle code understanding and implementation in code translation into two steps, formulating easier subproblems for LLMs to solve over complicated programs.
Some programs are compact, implementing multiple operations in a single dense statement by chaining several steps recursively or sequentially (Keshk and Dyer, 2023). Such programs may result from PL idioms or developers’ programming style (Keshk and Dyer, 2023). We observed that LLMs struggle to disentangle the functionality of such complicated programs and re-implement the functionality in the target PL via one-step direct translation. LLMs often miss or mess up steps in such cases. In comparison, a pseudocode-based transitive translation splits the code understanding of the complicated program and the implementation of the code intent and logic into two steps, reducing the burden on LLMs and promoting accuracy in each step.
Figure 6 illustrates an example of this case. Specifically, the original program includes only one Python statement that recursively conducts four operations (LABEL:lst:_rq1-type3-py2js-python). When translating this compact program via the one-step direct translation, even the powerful Qwen32B missed the row-sorting step in the output JavaScript program (LABEL:lst:_rq1-type3-py2js-direct). In comparison, when asked to describe the code intent and logic using pseudocode, Qwen32B was able to interpret the complicated logic in the compact Python program in pseudocode (LABEL:lst:_rq1-type3-py2js-pseudocode). The pseudocode then promotes the successful implementation of the operation JavaScript (LABEL:lst:_rq1-type3-py2js-viapseudocode). Similar symptoms are also observed on Phi-4 and Qwen7B.
@@
Listing 1: Source Python Program to Translate
@@
Listing 2: Pseudocode of Source Python Program
@@
Listing 3: Translation with Direct
@@
Listing 4: Translation with viaPseudocode (w/ code)
Figure 6. An Example of Python-to-JavaScript Translation Involving Compact Original Programs (Qwen32B)3
@@ 1. Pro-2
We observed that LLM-based direct translation may overlook the differences in the features of source and target PLs, resulting in incompatible implementations violating the mechanism of the target PL. Meanwhile, pseudocode describes the code’s intent and logic, hiding PL-specific implementation. When generating target code from the pseudocode, LLMs often properly implement the intended function in a manner fitting the mechanism of the target PL, and are less likely to be distracted by the incompatible patterns of source PL in the original program.
Footnote 2 shows an example of this case with a Go-to-Rust translation. Specifically, when translating a Go program with an inner function (LABEL:lst:_rq1-type1-go2rust-go) into Rust via direct translation, Qwen7B mechanically followed the Go program to declare a common inner function in Rust without realizing the difference in default closure capture modes in Go and Rust and forgetting to indicate a mutable closure capture mode (LABEL:lst:_rq1-type1-go2rust-direct). Meanwhile, based on pseudocode that mainly indicates the code intent and logic (LABEL:lst:_rq1-type1-go2rust-pseudocode), Qwen7B correctly implemented the functionality in a proper manner that fits Rust’s memory management (LABEL:lst:_rq1-type1-go2rust-viapseudocode). Interestingly, all five studied LLMs failed with the same fault in direct translation while succeeding with the help of pseudocode. This also echoes the helpfulness of pseudocode-based code translation in translating from flexible PLs to more rigid PLs observed in RQ2 (Section 4.2).
@@
Listing 5: Source Go Program to Translate
@@
Listing 6: Pseudocode of Source Go Program
@@
Listing 7: Translation with Direct
@@
Listing 8: Translation with viaPseudocode (w/ code)
Figure 7. An Example of Go-to-Rust Translation Involving Unique Closure Capturing in Go and Rust (Qwen7B)333“…” represents code less related to the illustration and omitted for saving space. The comment is not generated by LLMs but annotated by authors for case explanation.
@@ 1. Pro-3
We observed that LLMs could hallucinate incorrect implementations that fail to realize the functionality of the PL-specific APIs in original programs during direct translation. LLMs struggle to identify or implement a semantically equivalent API counterpart in the target PL based on only the original program, and may even hallucinate non-existing APIs. Meanwhile, pseudocode often describes the functionality implemented with PL-specific APIs in natural language, guiding LLMs to figure out appropriate APIs and API usages in the target PL as they work out a code generation task without being misled by concrete API usages in original programs.
Figure 8 illustrates an example of this case. Specifically, the original Python program uses an API “bit_count” to calculate the number of ones in the binary expression of an integer (LABEL:lst:_rq1-type2-py2go-python). Due to no “bit_count” API in Go, when translating the Python program into Go via direct translation, Phi-4 implemented bit counting with a series of bit operations (LABEL:lst:_rq1-type2-py2go-direct). However, the generated bit operations do not align with the functionality of “bit_count”. Similarly, GPT-4o-mini hallucinated incorrect bit operations, and DSCoder even hallucinated a non-existing API. Instead, the pseudocode generated by Phi-4 clarified the functionality implemented by “bit_count” in Python (LABEL:lst:_rq1-type2-py2go-pseudocode), which guided Phi-4 to generate a correct subroutine “countBits” in Go to realize the functionality of “bit_count” (LABEL:lst:_rq1-type2-py2go-viapseudocode). DSCoder and GPT-4o-mini also generated a correct translation based on the pseudocode generated by themselves.
@@
Listing 9: Source Python Program to Translate
@@
Listing 10: Pseudocode of Source Python Program
@@
// index := (k - 1).bit_count() //non-existing API
Listing 11: Translation with Direct
@@
count := 0
for n > 0 {
@@
Listing 12: Translation with viaPseudocode (w/ code)
Figure 8. An Example of Python-to-Go Translation Involving Non-Existing Equivalent APIs in Go (Phi-4)3
Typical limitations of pseudocode-based code translation. We also observed three major limitations of pseudocode that cause errors in semantic code translation results. They stem from LLMs’ incorrect understanding of original programs, the semantic loss in information transmission, and the ambiguity of natural language. Enhancements to generate more accurate pseudocode are needed to mitigate these issues and harness the advantages of pseudocode-based code translation.
@@ 1. Limit-1
Explicit semantic code translation based on pseudocode introduces a compulsory code understanding step during translation. However, LLMs are not free from erroneous code understanding. An incorrect code intent or logic can conversely disturb code translation. For example, translating a straightforward Python expression “sum = sum+nums[i] if nums[i]<=sum else nums[i]” into Java may not necessitate the understanding of the code functionality. An incorrect intent conversely misled translation: Qwen32B hallucinated a code intent of “sum = max(sum, sum + nums[i])” for this expression, misleading the generation of the Java program. The issue happened with both $\mathcal{P}$ and $\mathcal{PC}$ strategies, while direct translation ($\mathcal{D}$) leads to a correct translation.
@@ 1. Limit-2
LLM-generated pseudocode may also miss certain essential information for implementing the complete functionality of the original program. For example, the data type of variables is often omitted in the LLM-generated pseudocode without any description of precision concerns, leaving operations sensitive to data type (e.g., requiring certain precision) hard to reproduce. Another typical missing information for reproducing code semantics is the logic of complicated algorithms. From a concise description without detailed implementations, LLMs may not accurately reproduce all algorithms. The concise description may also miss customized operations in algorithms, leaving the generated code degenerate into a general implementation lacking the expected features.
@@ 1. Limit-3
Although natural language can depict PL-agnostic code intent and logic, its nature of ambiguity may introduce noise and mislead the code generation step. A typical issue observed in the failure cases is the ambiguity of “to” and “downto” in “for” loops. Such expressions cannot clarify the inclusion of the border value in the loop. The inclusion of the original program (i.e., using $\mathcal{PC}$ strategy) cannot always mitigate the disturbance caused by the ambiguity as well. In comparison, a direct translation from the original program does not suffer from such ambiguity-induced issues.
Research Opportunities. To leverage the advantages of pseudocode-based code translation observed in the previous three RQs and its potential revealed in RQ4, we suggest four directions for future enhancement based on our findings. These target the generation and validation of pseudocode, as well as hybrid approaches to combine the strengths of direct and transitive translation.
@@ 1. Chance-2
Effective validation strategies may help identify inconsistencies and semantic losses. For example, round-trip validation (Aiken and Park, 2010; Zhou and Sun, 2018) may identify buggy pseudocode by translating the generation result back to the source PL, and LLMs may infer the intended behavior based on code (Xu et al., 2025) and compare pseudocode to it to debug pseudocode. Incorporating such validation loops would help identify and mitigate failures stemming from low-quality pseudocode suffering from the three limitations.
@@ 5. Threats to Validity
Table 4. Pass@10 Rates on All 323 Tasks and 114 Newer Tasks After Cut-off Date (2024-06-01)
Strategy LLM All 323 Tasks
114 Newer Tasks
LLM All 323 Tasks
114 Newer Tasks
$\mathcal{D}$(Direct Translation) Qwen 32B 0.8777 0.8604 Qwen 7B 0.7934 0.7623
$\mathcal{P}$(viaPseudocode) 0.8703 (-0.84%) 0.8424 (-2.09%) 0.7446 (-21.43%) 0.5320 (-30.21%)
$\mathcal{PC}$(viaPseudocode w/ prog.) 0.9023 (+2.80%) 0.8848 (+2.84%) 0.8016 (+2.96%) 0.7435 (-2.47%)
$\mathcal{D}\&\mathcal{P}$(Hybrid [DT&vP]) 0.9325 (+6.24%) 0.9199 (+6.92%) 0.8587 (+4.92%) 0.7956 (+4.37%)
$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.9292 (+5.87%) 0.9170 (+6.58%) 0.8616 (+6.93%) 0.8151 (+6.93%)
@@
$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.8886 (+6.25%) 0.8735 (+7.23%) 0.8616 (+8.60%) 0.8359 (+8.95%)
$\mathcal{D}$(Direct Translation) GPT-4o -mini 0.8539 0.8363 Avg 0.8309 0.8082
$\mathcal{P}$(viaPseudocode) 0.7409 (-13.23%) 0.7155 (-14.44%) 0.7769 (-9.04%) 0.7157 (-11.44%)
$\mathcal{PC}$(viaPseudocode w/ prog.) 0.8590 (+0.60%) 0.8520 (+1.88%) 0.8409 (+2.48%) 0.8161 (+0.98%)
$\mathcal{D}\&\mathcal{P}$(Hybrid [DT&vP]) 0.9031 (+5.76%) 0.8857 (+5.91%) 0.8891 (+6.37%) 0.8623 (+6.70%)
$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.9093 (+6.49%) 0.8962 (+7.16%) 0.8901 (+6.80%) 0.8675 (+7.35%)
Representativeness of Studied Subjects. The first threat to our study is about the representativeness and generalizability of the observations on our study subjects. To mitigate this threat, we followed existing work to investigate code translation across six PLs. These include Python, C++, Java, and JavaScript, which are widely adopted in daily development and studied in existing works (TIOBE Index, [n. d.]; Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), as well as the emerging Go and Rust (Pan et al., 2024; Macedo et al., 2025). Our study results are expected to guide the adoption of pseudocode-based translation for these popular PLs.
We constructed translation tasks based on LeetCode problems in LiveCodeBench (Jain et al., 2025). LeetCode problems are widely used in code translation studies (Wang et al., 2023a; Ahmad et al., 2023) and benchmark of LLMs’ coding ability (Hui et al., 2024; Team, 2024; Jain et al., 2025) since they provide adequate test cases to evaluate candidate solutions and have easy-to-collect multilingual solution programs. In addition, these tasks require common engineering and algorithm implementations useful in daily development. Thus, we consider the study results on their solution programs should be meaningful and generalizable to practical development code.
We used five LLMs deployable on our machine as the code translator. These include both general LLMs and coder LLMs, and both closed-source and open-source LLMs from four families. They rank a varying range (from Top-10 to 68) on the famous BigCodeBench (Zhuo et al., 2025). Also, the training data for these LLMs has a cutoff date with relatively little overlap with the timeline of LeetCode problems. Thus, we consider the results learned with these LLMs should be representable and reliable.
Data Contamination. A potential threat to our study result is the subject to the data contamination issue, where LLMs perform well because they have learned the ground truth during training rather than pseudocode-based translation strategies. Although our translation tasks based on LiveCodeBench are much newer than the subjects in the conventional code translation benchmarks as introduced in Section 3.2, there is still some overlap between the timeline of LiveCodeBench problems and the studied LLMs’ training data. To further verify the cause of the observed performance improvement, we re-evaluate LLMs on the solution programs of a clean subset of 114 LeetCode problems released after the latest knowledge cutoff of the studied LLMs, i.e., June 1, 2024. As shown in Table 4, the performance improvement of different strategies is generally consistent with that observed on all 323 tasks. Thus, we consider the observation based on the complete set is meaningful and provides more statistically meaningful observations with more subjects.
@@
Representativeness of Studied Translation Strategies. The representativeness of the studied translation strategies and prompts is another potential threat to our study, impacting the meaningfulness of our observations. To mitigate this threat, we studied five translation strategies, including one direct translation strategy, which is the commonly used approach in existing studies and practices, as well as two pseudocode-based translation strategies and two hybrid translation strategies. The pseudocode-based strategies emulate the human practices of semantic translation on natural languages (Newmark, 1981) and are implemented in a transitive translation manner following (Macedo et al., 2025). The hybrid strategies are based on straightforward result combination to explore the helpfulness of the combined advantages of strategies, considering the feasibility of selecting multiple translation candidates based on original programs (Macedo et al., 2025). We carefully designed the prompts following existing studies of direct code translation (Yang et al., 2024; Pan et al., 2024) and pseudocode-based code generation (Wu et al., 2025). Thus, they can reflect the typical usage of these strategies and prompts and lead to meaningful observations.

### v3

@@ 4.5. Discussion: Advantages and Limitations of Pseudocode-Based Translation
After understanding the effectiveness of pseudocode-based code translation from the quantitative results in four RQs, we further conduct case studies to attribute its successes and failures to learn concrete insights for the effective adoption of this approach. Specifically, two authors with over five years of programming experience analyzed the results of 200 randomly sampled translation tasks (i.e., 100 tasks where direct translation failed while pseudocode-based translation succeeded, and another 100 tasks where direct translation succeeded while pseudocode-based translation failed [^3] To study reasons of clear successes/failures rather than randomness, we only analyzed the tasks where over five in ten attempts of a strategy passed all tests (successful tasks) and where none attempt passed all tests (failed tasks).), to identify the advantages/limitations of pseudocode for successful/failed translations compared with direct translation. These tasks are sampled across all the studied LLMs and PL pairs, representing a 95% confidence level with a 10% margin of error. The participants first discussed potential categorizations and criteria of advantages and limitations via a pilot study on 40 sampled tasks. Then, they independently annotated each task (with substantial agreement indicated by Cohen’s Kappa coefficients (Cohen, 1960) over 0.78), and next discussed to reach a consensus on all the annotations. Each task can be annotated with multiple advantages/limitations of pseudocode. In addition, based on all the findings in this work, we further discuss several research directions to mitigate the limitations and harness the potential for pseudocode-based code translation.
4.5.1. Key Advantages of Pseudocode-based Code Translation.
We identified three major advantages of pseudocode-based translation over direct translation. One of them facilitates the understanding of the original programs and reduces the burden of an integrated step, and the other two help LLMs handle differences across PLs. They align with the limitations of direct translation identified in prior studies (Pan et al., 2024; Jiao et al., 2023; Macedo et al., 2025). Practitioners are encouraged to leverage these advantages to translate the original programs or across PL pairs fitting them.
@@ 1. Pro-1
    Pseudocode helps decouple code understanding and implementation in code translation, formulating easier subproblems for LLMs to solve when translating complicated programs.
Some programs are compact, implementing multiple operations in a single dense statement by chaining several steps recursively or sequentially. Such programs may result from PL idioms or developers’ programming style (Keshk and Dyer, 2023). We observed that LLMs struggle to understand the tangled functionality of such complicated programs and re-implement the functionality in the target PL via one-step direct translation. LLMs often miss or mess up steps in such cases. In comparison, a pseudocode-based transitive translation splits the code understanding of the complicated program and the implementation of the code intent and logic into two steps, reducing the burden on LLMs and promoting accuracy in each step. This advantage was observed in 34% of analyzed cases where pseudocode-based translation succeeded while direct translation failed.
Figure 7 illustrates an example of this case. Specifically, the original Python program includes only one statement that recursively conducts four operations (LABEL:lst:_rq1-type3-py2js-python). When translating this compact program via the one-step direct translation, even the powerful Qwen32B missed the row-sorting step in the output JavaScript program (LABEL:lst:_rq1-type3-py2js-direct). In comparison, when asked to describe the code intent and logic using pseudocode, Qwen32B was able to interpret the complicated logic of the compact Python code using pseudocode (LABEL:lst:_rq1-type3-py2js-pseudocode). The pseudocode then promotes the successful implementation of the operation in JavaScript (LABEL:lst:_rq1-type3-py2js-viapseudocode). Phi-4 and Qwen7B also suffered from similar issues when conducting direct translation and succeeded in pseudocode-based translation.
@@
Listing 1: Original Python Program to Translate
@@
Listing 2: Pseudocode of Original Python Program
@@
Listing 3: Direct Translation Result
@@
Listing 4: Pseudocode-based Translation Result
Figure 7. An Example of Python-to-JavaScript Translation Involving Compact Original Programs (Qwen32B)5
@@ 1. Pro-2
We observed that LLM-based direct translation may overlook the differences in the features of source and target PLs, resulting in incompatible implementations in the translation result that violate the mechanism of the target PL. Meanwhile, pseudocode describes the original program’s intent and logic, hiding PL-specific implementation. When generating target code from pseudocode, LLMs often properly implement the intended function in a manner fitting the mechanism of the target PL, and are less likely to be distracted by the incompatible patterns of source PL in original programs. This advantage was observed in 56% of analyzed cases where pseudocode-based translation outperformed direct translation.
Footnote 4 shows an example of this case with a Go-to-Rust translation. Specifically, when translating a Go program with an inner function (LABEL:lst:_rq1-type1-go2rust-go) into Rust via direct translation, Qwen7B mechanically followed the Go program to declare a common inner function in Rust without realizing the difference in default closure capture modes in Go and Rust and failing to indicate a mutable closure capture mode in Rust (LABEL:lst:_rq1-type1-go2rust-direct). Meanwhile, based on pseudocode that mainly indicates the code intent and logic (LABEL:lst:_rq1-type1-go2rust-pseudocode), Qwen7B correctly implemented the functionality in a proper manner that fits the memory management of Rust (LABEL:lst:_rq1-type1-go2rust-viapseudocode). Notably, all five studied LLMs failed with the same fault in direct translation while succeeding with the help of pseudocode. This also echoes the helpfulness of pseudocode-based code translation in translating from flexible PLs to more rigid PLs observed in RQ2 (Section 4.2).
@@
Listing 5: Original Go Program to Translate
@@
Listing 6: Pseudocode of Original Go Program
@@
Listing 7: Direct Translation Result
@@
Listing 8: Pseudocode-based Translation Result
Figure 8. An Example of Go-to-Rust Translation Involving Unique Closure Capturing in Go and Rust (Qwen7B)555“…” represents code less related to the illustration and omitted for saving space. The comment is not generated by LLMs but annotated by authors for case explanation.
@@ 1. Pro-3
We observed that LLMs could hallucinate incorrect implementations that fail to realize the functionality of the PL-specific APIs in original programs during direct translation. LLMs struggle to identify or implement a semantically equivalent API counterpart in the target PL based on only the original program, and may even hallucinate non-existent APIs. Meanwhile, pseudocode often describes the functionality implemented with PL-specific APIs in natural language, guiding LLMs to figure out appropriate APIs and API usages in the target PL as they work out a code generation task without being misled by concrete API usages in original programs. This advantage was observed in 33% of analyzed cases where pseudocode-based translation outperformed direct translation.
Figure 9 illustrates an example of this case. Specifically, the original Python program uses an API “bit_count” to calculate the number of ones in the binary expression of an integer (LABEL:lst:_rq1-type2-py2go-python). Due to no “bit_count” API in Go, when translating the Python program into Go via direct translation, Phi-4 implemented bit counting with a series of bit operations (LABEL:lst:_rq1-type2-py2go-direct). However, the generated bit operations do not align with the functionality of “bit_count”. Similarly, GPT-4o-mini hallucinated incorrect bit operations, and DSCoder even hallucinated a non-existent API. Instead, the pseudocode generated by Phi-4 clarified the functionality implemented by “bit_count” in Python (LABEL:lst:_rq1-type2-py2go-pseudocode), which guided Phi-4 to generate a correct subroutine “countBits” in Go to realize the functionality of “bit_count” (LABEL:lst:_rq1-type2-py2go-viapseudocode). DSCoder and GPT-4o-mini also generated a correct translation based on the pseudocode generated by themselves.
@@
Listing 9: Original Python Program to Translate
@@
Listing 10: Pseudocode of Original Python Program
@@
// index := (k - 1).bit_count() //non-existent API
Listing 11: Direct Translation Result
@@
count := 0 // Counting bits in the integer as
for n > 0 { // the Python API bit_count does.
@@
Listing 12: Pseudocode-based Translation Result
Figure 9. An Example of Python-to-Go Translation Involving Non-Existent Equivalent APIs in Go (Phi-4)5
4.5.2. Typical Limitations of Pseudocode-based Code Translation.
We also observed three major limitations that cause errors in pseudocode-based code translation due to issues in the generated pseudocode. They stem from LLMs’ incorrect understanding of original programs, the semantic loss in information transmission, and the ambiguity of natural language. Enhancements to generate more accurate pseudocode are needed to mitigate these limitations and harness the advantages of pseudocode-based code translation.
@@ 1. Limit-1
Explicit semantic code translation based on pseudocode introduces a compulsory code understanding step during translation. However, LLMs are not free from erroneous code understanding. An incorrect code intent or logic can conversely disturb code translation. For example, translating a straightforward Python expression “sum = sum+nums[i] if nums[i]<=sum else nums[i]” into Java may not necessitate the understanding of the code functionality. An incorrect intent conversely misled translation: Qwen32B hallucinated a code intent of “sum = max(sum, sum + nums[i])” for this expression in pseudocode, misleading the generation of the Java program. The issue happened with both $\mathcal{P}$ and $\mathcal{PC}$ strategies, while direct translation ($\mathcal{D}$) leads to a correct translation. This limitation was observed in 66% of analyzed cases where pseudocode-based translation failed while direct translation succeeded.
@@ 1. Limit-2
LLM-generated pseudocode may also miss certain essential information for implementing the complete functionality of the original program. For example, the data type of variables is often omitted in the LLM-generated pseudocode without any description of precision concerns, leaving operations sensitive to data type (e.g., requiring certain precision) hard to reproduce. Another typical missing information for reproducing code semantics is the logic of complicated algorithms. From a brief description of logic without detailed implementations, LLMs may not accurately reproduce all algorithms. The concise description may also miss customized operations in algorithms, leaving the generated code degenerate into a general implementation lacking the expected features. This limitation was observed in 27% of analyzed cases where direct translation outperformed pseudocode-based translation.
@@ 1. Limit-3
Although natural language can depict PL-agnostic code intent and logic, its nature of ambiguity may introduce noise and mislead the code generation step. A typical ambiguity is about the descriptions of loop boundaries using “to”/“downto”, which cannot clarify the inclusion of border values. For example, Qwen32B generated a pseudocode with “for i from 0 to length of nums” for the original Rust loop “for i in 0..nums.len()”, where the upper border value is excluded. When implementing the pseudocode in Go, Qwen32B generated a loop “for i:=0; i<=len(nums) …” with the upper border value included, resulting in a semantic drift. Another common issue is the unclear precedence among arithmetic operations described by natural language descriptions. For example, Qwen32B generated a pseudocode using “a + integer division of (k - 1) by a - 1” to describe the original Java expression “a+(k-1)/a-1”. The unclear precedence of addition and subtraction misled Qwen32B to implement “a+Math.floor((k-1)/(a-1))” in JavaScript, which drifts from original semantics. Including original programs (i.e., using $\mathcal{PC}$ strategy) cannot mitigate the disturbance caused by the ambiguity as well. In comparison, a direct translation from the original program does not suffer from such ambiguity-induced issues. This limitation was observed in 17% of analyzed cases where direct translation outperformed pseudocode-based translation.
4.5.3. Research Opportunities.
To leverage the advantages of pseudocode-based code translation observed in the previous three RQs and its potential revealed in RQ4, we suggest four directions for future enhancement based on our findings. These target the generation and validation of pseudocode, as well as hybrid approaches to combine the strengths of direct and transitive translation.
@@ 1. Chance-2
Effective validation strategies may help identify inconsistencies and semantic losses. For example, round-trip validation (Aiken and Park, 2010; Zhou and Sun, 2018) may help identify buggy pseudocode by translating the generation result back to the source PL for comparison, and LLMs may infer the intended behavior based on code (Xu et al., 2025) and compare pseudocode to it to debug pseudocode. Incorporating such validation steps in the loop would help identify and mitigate failures stemming from low-quality pseudocode suffering from the three limitations.
@@ 1. Chance-4
  1. Chance-5
    Enhancing code translation via reasoning.
Several advanced LLMs (e.g., DeepSeek-R1) have a built-in mechanism to generate reasoning traces in the response process, which proves helpful in guiding LLMs to solve complicated tasks like mathematical and programming problems (Guo et al., 2025). However, it remains underexplored how to effectively leverage reasoning traces for code translation. There can be several potential directions. One idea is to directly use reasoning traces as an intermediate representation to guide code translation, similar to pseudocode. However, reasoning traces are often more verbose and less structured than pseudocode, which may introduce unique challenges. For example, we observed serious overthinking issues when prompting DeepSeek-R1 Distilled Qwen7B to generate reasoning traces before code translation, leading to a decrease in translation accuracy compared to direct and pseudocode-based translation and an increase in computation cost, like issues revealed by existing studies (Chen et al., 2025b). In addition, the adoption of reasoning during pseudocode generation is also worth exploring to help LLMs better understand and depict original programs. In this study, we have employed chain-of-thought prompting (Wei et al., 2022) to trigger reasoning traces of the five LLMs studied, aiming to guide LLMs to explicitly put down the abstraction and analysis steps that may help generate correct and concise pseudocode. Nevertheless, more systematic designs to effectively integrate reasoning into code translation are worth further exploration.
@@ 5. Threats to Validity
Representativeness of Studied Subjects. The first threat to our study is about the representativeness and generalizability of the observations on our study subjects. To mitigate this threat, we followed existing work to investigate code translation across six PLs. These include Python, C++, Java, and JavaScript, which are widely adopted in daily development and studied in existing works (TIOBE Index, ; Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), as well as the emerging Go and Rust (Pan et al., 2024; Macedo et al., 2025). Our study results are expected to guide the adoption of pseudocode-based translation for these popular PLs.
We constructed translation tasks based on various LeetCode problems in LiveCodeBench (Jain et al., 2025). LeetCode problems are widely used in code translation studies (Wang et al., 2023a; Ahmad et al., 2023) and benchmarking LLMs’ coding ability (Hui et al., 2024; Team, 2024; Jain et al., 2025) since they provide high-quality tests to evaluate candidate solutions and provide easy access to collect multilingual programs. Also, they require common engineering and algorithm implementations useful in daily development. Thus, we consider the study results on their solution programs should be meaningful and generalizable to practical development code.
We used five LLMs deployable on our machine as the code translator, including both general and coding LLMs, and both closed-source and open-source LLMs from four families. They rank a varying range (from Top-10 to 68 according to the record on Oct 1, 2025) on the famous BigCodeBench (Zhuo et al., 2025). Also, the training data for these LLMs has a cutoff date with relatively little overlap with the timeline of the LeetCode problems to construct translation tasks. Therefore, we consider the results learned with these LLMs should be representable and reliable.
Data Contamination. A potential threat to our study result is the subject to the data contamination issue, where LLMs perform well because they have learned the ground truth during training, rather than pseudocode-based translation strategies. Although our translation tasks built upon LiveCodeBench are much newer than the subjects in the conventional code translation benchmarks as introduced in Section 3.2, there is still some overlap between the timeline of LiveCodeBench problems and the studied LLMs’ training data. To further verify the cause of the observed performance improvement, we re-evaluate LLMs on the solution programs of a clean subset of 114 problems released after the latest knowledge cutoff of the studied LLMs, i.e., June 1, 2024. As shown in Table 6, the performance improvement of different strategies is generally consistent with that observed on all 323 problems. Therefore, we consider the observation based on the complete set to be meaningful and provide more statistically meaningful observations with more subjects.
Table 6. Pass@10 Rates on All 323 Problems and 114 Newer Problems After Cut-off Date (2024-06-01)
Strategy LLM All 323 Probs
114 Newer Probs
LLM All 323 Probs
114 Newer Probs
$\mathcal{D}$(Direct Translation) Qwen 32B 0.8777 0.8604 Qwen 7B 0.7971 0.7623
$\mathcal{P}$(viaPseudocode) 0.8703 (-0.84%) 0.8424 (-2.09%) 0.6193 (-22.31%) 0.5320 (-30.21%)
$\mathcal{PC}$(viaPseudocode w/ ctx.) 0.9023 (+2.80%) 0.8848 (+2.84%) 0.7954 (-0.21%) 0.7435 (-2.47%)
$\mathcal{D}\&\mathcal{P}$(Hybrid[DT&vP]) 0.9325 (+6.24%) 0.9199 (+6.92%) 0.8363 (+4.92%) 0.7956 (+4.37%)
$\mathcal{D}\&\mathcal{PC}$(Hybrid[DT&vPw/c.]) 0.9292 (+5.87%) 0.9170 (+6.58%) 0.8523 (+6.93%) 0.8151 (+6.93%)
@@
$\mathcal{D}\&\mathcal{PC}$(Hybrid[DT&vPw/c.]) 0.8886 (+6.25%) 0.8735 (+7.23%) 0.8616 (+8.60%) 0.8359 (+8.95%)
$\mathcal{D}$(Direct Translation) GPT-4o -mini 0.8539 0.8363 (Avg) 0.8317 0.8082
$\mathcal{P}$(viaPseudocode) 0.7409 (-13.23%) 0.7155 (-14.44%) 0.7518 (-9.60%) 0.7157 (-11.44%)
$\mathcal{PC}$(viaPseudocode w/ ctx.) 0.8590 (+0.60%) 0.8520 (+1.88%) 0.8396 (+0.95%) 0.8161 (+0.98%)
$\mathcal{D}\&\mathcal{P}$(Hybrid[DT&vP]) 0.9031 (+5.76%) 0.8857 (+5.91%) 0.8846 (+6.37%) 0.8623 (+6.70%)
$\mathcal{D}\&\mathcal{PC}$(Hybrid[DT&vPw/c.]) 0.9093 (+6.49%) 0.8962 (+7.16%) 0.8882 (+6.80%) 0.8675 (+7.35%)
@@
Representativeness of Studied Translation Strategies. The representativeness of the studied translation strategies and prompts is another potential threat to our study, affecting the meaningfulness of our observations. To mitigate this threat, we studied five code translation strategies, including one direct translation strategy commonly used in LLM-based code translation studies and practices (Pan et al., 2024; Yang et al., 2024), as well as two pseudocode-based translation strategies and two hybrid translation strategies. The pseudocode-based strategies emulate the human practices of semantic translation on natural languages (Newmark, 1981) and are implemented in a transitive translation manner following (Macedo et al., 2025). The hybrid strategies are based on straightforward result combination to explore the helpfulness of the combined advantages of strategies, considering the feasibility of selecting multiple translation candidates based on original programs (Macedo et al., 2025). We carefully designed the prompts following existing studies of direct code translation (Yang et al., 2024; Pan et al., 2024) and pseudocode-based code generation (Wu et al., 2025). Thus, they can reflect the typical usage of these strategies and prompts and lead to meaningful observations.